The inverse of the demand and supply functions for shoes is given by the

following equations respectively:

Demand: P = 1400 - 2Qd

Supply: P = 200 + 1Qs

1.1. Calculate the equilibrium price and quantity of shoes.

(5)

1.2. Assume that the price of shoes is R700. Use your answer in 1.1 to explain

the resulting situation in the market for shoes, and how equilibrium will

be restored without government intervention, ceteris paribus


1
Expert's answer
2021-11-15T17:52:26-0500

Solution:

1.1.). At equilibrium: Qd = Qs

1400 – 2Q = 200 + 1Q

1400 – 200 = 1Q + 2Q

1200 = 3Q

Q = 400

Equilibrium quantity = 400

Substitute in either the demand or supply function to derive equilibrium price:

P = 1400 – 2Qd = 1400 – 2(400) = 1400 – 800 = 600

Equilibrium price = 600

 

1.2). The price of R700 will be above the equilibrium price, which means that the quantity of shoes supplied will exceed the quantity of shoes demanded, resulting in shoes surplus in the market.